# Scalable LLM Math Reasoning Acceleration with Low-rank Distillation

### v2

@@ 4 Experiments
We showcase the effectiveness of Caprese at recovering much, if not all, of the math performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks. In the following section (Section 4.1), we observe better math performance for instruct LLMs without losing their generalizability. Then, we study Caprese’s benefit on two axes of test-time scaling: number of generations (Section 4.2) and generation length with thinking models (Section 4.3). Finally, we highlight Caprese’s latency and length improvements in Section 4.4. When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying sparse method and similarly for GRIFFIN. Otherwise, we use “Caprese” for brevity. Unless specified, the FF intermediate feature sparsity is set at 50%. NOTE: For a more meaningful baseline, we only apply GRIFFIN to the first half of the model for all experiments since we observed much steeper drops in math performance if used for all layers. CATS is applied to all layers.
4.1 Instruct Models
We first look into different instruct LLMs which are performative on non-reasoning tasks and have decent reasoning performance. We show Caprese is able to preserve math performance without sacrificing quality on pure language tasks like question answering. We test on the Llama 3 family [DJP+24] and Gemma 2 family [TRP+24] of models on 0-shot GSM8K [CKB+21], MATH [HBK+21], CoQA [RCM19], and QASPER [DLB+21], using CoT prompts for math tasks.
Table 1 shows that Caprese is able to preserve most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math. In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese is able to effectively recover the lost performance. E2E Caprese is the most performative in the majority of math scenarios, suggesting a benefit beyond minimizing local error. Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A). The best performers for CoQA and QASPER are a toss-up, but all methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases). CATS and Caprese considerably improve GSM8K performance of Llama 3.1 8B Instruct, which has trouble following the prompt template with zero-shot.
4.2 Scaling Best-of-$N$: Coverage
Now, we see the generalizability of Caprese on the first axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$. Increasing the number of generations per prompt, $N$, is one way to scale inference compute for improved performance, as the probability of generating a correct answer increases. To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH. We use an oracle verifier to accurately assess the quality of the pool of generated responses. To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample. Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3. Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
@@ 4.3 Scaling Length: Thinking Models
Model MATH-500 AIME 2024 AMC 2023 BRUMO 2025
DS-Qwen 1.5B 79.40 30.00 $\pm$ 2.72 60.83 $\pm$ 7.47 25.83 $\pm$ 3.19
GRIFFIN 42.00 2.50 $\pm$ 1.67 16.88 $\pm$ 1.25 1.67 $\pm$ 1.92
Layer-wise Caprese 47.20 2.50 $\pm$ 1.67 28.13 $\pm$ 2.39 0.00 $\pm$ 0.00
E2E Caprese 60.40 6.67 $\pm$ 2.72 34.38 $\pm$ 6.57 4.17 $\pm$ 1.67
CATS 72.00 11.67 $\pm$ 6.94 37.50 $\pm$ 3.54 10.83 $\pm$ 4.19
Layer-wise Caprese 73.80 15.83 $\pm$ 3.19 48.13 $\pm$ 3.75 13.33 $\pm$ 2.72
E2E Caprese 74.80 20.83 $\pm$ 1.67 47.50 $\pm$ 3.54 20.00 $\pm$ 2.72
DS-Qwen 7B 90.20 47.50 $\pm$ 7.87 75.00 $\pm$ 4.08 45.83 $\pm$ 4.19
GRIFFIN 80.60 21.67 $\pm$ 4.30 62.50 $\pm$ 8.90 25.00 $\pm$ 8.39
Layer-wise Caprese 84.80 29.17 $\pm$ 3.19 64.38 $\pm$ 7.18 27.50 $\pm$ 4.19
E2E Caprese 85.40 30.00 $\pm$ 6.09 71.88 $\pm$ 5.54 29.17 $\pm$ 5.69
CATS 89.20 34.17 $\pm$ 7.88 62.50 $\pm$ 2.04 37.50 $\pm$ 7.39
Layer-wise Caprese 87.00 35.00 $\pm$ 4.30 70.00 $\pm$ 10.21 39.17 $\pm$ 3.19
E2E Caprese 90.00 33.33 $\pm$ 4.71 78.13 $\pm$ 5.15 47.50 $\pm$ 5.00
DS-Qwen 14B 92.80 63.33 $\pm$ 2.72 90.63 $\pm$ 2.39 60.83 $\pm$ 4.19
GRIFFIN 89.80 32.50 $\pm$ 7.39 80.63 $\pm$ 3.75 33.33 $\pm$ 3.85
Layer-wise Caprese 90.80 41.67 $\pm$ 5.77 86.88 $\pm$ 4.27 43.33 $\pm$ 4.71
E2E Caprese 89.20 39.17 $\pm$ 9.18 84.38 $\pm$ 6.57 40.00 $\pm$ 2.72
CATS 92.80 58.34 $\pm$ 1.92 88.75 $\pm$ 1.44 55.00 $\pm$ 3.33
Layer-wise Caprese 91.00 61.67 $\pm$ 5.77 88.75 $\pm$ 1.44 53.00 $\pm$ 7.20
E2E Caprese 92.00 58.34 $\pm$ 9.62 87.50 $\pm$ 2.04 49.17 $\pm$ 4.19
Table 2: Thinking models’ 0-shot accuracies on math reasoning tasks. Sparsity is set at 50%, and $r=256$. AIME 2024 and AMC 2023 columns also include the sample standard deviation across 4 runs. Further improvements with reselection are shown in Table 3.
Thinking models provide another axis of test-time scaling by augmenting the CoT length before giving a final answer. Because this entails very long generation lengths, it is critical that error accumulation across tokens be minimized. We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [YYZ+24, GYZ+25] using the prompt templates and configurations from Open R1 [Fac25] (temperature and top-$p$ set to 0.6 and 0.95, respectively). For these models, we test 0-shot performance on MATH-500 [HBK+21, LKB+23], AIME 2024 [AIM25], AMC 2023 [AMC23], and BRUMO 2025 [BRU25] for a max generation length of 32768. Due to the small size of AIME 2024, AMC 2023, and BRUMO 2025, we evaluate them 4 times, reporting the average accuracy and sample standard deviation.
From Table 2, Caprese is the most performative in most cases. DS-Qwen 1.5B and 7B see the greatest benefit with E2E Caprese usually achieving the highest accuracy, sometimes even exceeding the full model’s performance. DS-Qwen 7B with CATS brings AMC 2023 accuracy down to 62.50% from 75.00%, but Caprese lifts performance above the original model’s to 78.13%, and similarly with DS-Qwen 7B with CATS and BRUMO 2025. All methods are more robust as the model size increases. For instance, CATS and GRIFFIN had very little impact on DS-Qwen 14B’s accuracy on MATH-500, with Caprese performing similarly. AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with several degradation and partial recovery using our method. In the next section, we show a simple way to push this recovery even further with reselection.
@@ 4.3.1 Enhanced Performance with Reselection
For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$. While this can be done by re-running prefill on all tokens, there is a more efficient way by passing in the generated tokens following the last reselection through the full model. As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs. This is like verification in speculative decoding [CBI+23, LKM23]. As $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use (as described in Section 2). This updates the pruned layers yet avoids prefill for all tokens. Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease the number of decode steps between reselection rounds. Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, paired with Caprese and reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
@@
Reselection is also possible with CATS but requires recomputing prefill for all tokens. The parameter to update is the hard thresholding parameter $\tau$, but because this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory. In turn, prefill will need to be redone anytime we want to update $\tau$. Since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations. Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
@@ 4.4 Efficiency
Caprese reduces generation latency. Table 4 shows the latencies of different generation settings. Caprese cuts total latency and time to next token by >16% for the Qwen 2.5 14B architecture. Moreover, the latency differences between GRIFFIN and Caprese are exceptionally small, suggesting that Caprese has minimal overhead. We also include latencies of GRIFFIN paired with LoRA weights, which although faster than the base model, negates >40% of the time savings that GRIFFIN has provided. This highlights the efficiency suboptimality of LoRA during inference. Metrics were collected on an NVIDIA L40 GPU using BF16 precision.
@@
Table 4: End-to-end generation latency (s) and average time to next token (ms) for GRIFFIN, LoRA, and Caprese using Qwen 2.5 14B. For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
@@ 4.4.1 Bonus: Effect on Natural Response Length
Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model. Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference (except for cutting off generation at 32K tokens). Shown in Figure 4 with MATH-500, the shortest mean response length for all problem difficulties is either the full model or Caprese. Meanwhile, CATS consistently outputs the longest responses, averaging roughly 1K more across every sample. It is also interesting to note that with increasing problem difficulty, all models and methods naturally allocate more inference tokens towards answering the question. Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DS-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit. This brevity of Caprese is more pronounced in larger models, proportionally, from a 5.8% reduction in tokens compared to the full DS-Qwen 7B model to an 8.5% reduction compared to the full DS-Qwen 14B model. The cause of this phenomenon is of interest for future work. See Appendix D for similar observations with GRIFFIN.
@@ 5 Conclusion
To combat the inefficiency of the long and brittle generation process associated with math reasoning, we introduce Caprese, a highly performant and efficient method with strong math and language capabilities that is compatible with a broad class of efficient FF algorithms. In the future, it would be interesting to evaluate its benefit in other reasoning domains besides math, explore the use of input-dependent low-rank layers, and investigate the knowledge contained in these low-rank layers. Caprese, along with future developments, pushes towards the long-term goal of degradation-free efficient LLM inference.

### v3

@@ 4 Experiments
We showcase the effectiveness of Caprese at recovering much, if not all, of the reasoning performance lost from efficient inference algorithms without sacrificing efficiency or performance on language tasks. Beginning with sequence length scaling in Section 4.1, Caprese is able to significantly or completely recover lost reasoning performance from existing efficient algorithms in thinking LLMs. Next, we demonstrate Caprese’s benefit when scaling the number of generations, another axis of tes-time scaling, in Section 4.2. Then in Section 4.3, we investigate typical instruct LLM settings without test-time scaling where we again observe better math performance without losing their generalizability. Finally, we highlight Caprese’s latency and length improvements in Section 4.4. Although we focus on coupling Caprese with sparse FF methods, we show our method’s merit carries over to quantization in Appendix D, broadening the potential applications of Caprese. When ambiguous, we denote Caprese (CATS) to be our method with CATS as the underlying compression method and similarly for GRIFFIN. Otherwise, we use “Caprese” for brevity. Unless specified, FF intermediate feature sparsity is set at 50%. For a more meaningful baseline, we only apply GRIFFIN to the first half of models as we observe steeper drops in math performance if used for all layers. CATS is applied to all layers.
@@ 4.1 Scaling Length: Thinking Models
Thinking models scale inference by augmenting the CoT length before giving a final answer. Since this entails long generation lengths, it is critical that error accumulation across tokens be minimized. We test on DeepSeek-R1-Distill-Qwen (which we abbreviate to DS-Qwen) models [57, 23] using Open R1 [20] prompt templates and configurations (temperature and top-$p$ set to 0.6 and 0.95, respectively). We test 0-shot performance on math and science tasks MATH-500 [24, 34], AIME 2024 [1], AMC 2023 [2], BRUMO 2025 [6], and GPQA [44], for a max generation length of 32768. Due to the small size of AIME 2024, AMC 2023, and BRUMO 2025, we evaluate them 4 times, reporting the average accuracy and sample standard deviation. We observe none of the methods have any significant effect on standard deviation in Table 2.
From Table 1, Caprese is the most performative in most cases. DS-Qwen 1.5B and 7B see the greatest benefit with E2E Caprese usually achieving the highest accuracy, sometimes even exceeding the full model’s performance. DS-Qwen 7B with CATS brings AMC 2023 accuracy down to 62.50% from 75.00%, but Caprese lifts performance above the original model’s to 78.13%, and similarly with DS-Qwen 7B with CATS and BRUMO 2025. All methods are more robust as model size increases. AIME 2024 and BRUMO 2025 challenge CATS and GRIFFIN with severe degradation and partial recovery with Caprese. In the next section, we show a simple way to push this recovery even further with reselection.
Table 2: Average thinking model performance standard deviation on AIME 2024, AIME 2023, and BRUMO 2025 from Table 1. There is no clear consistent relationship between Caprese and accuracy variance.
DS-Qwen 1.5B DS-Qwen 7B DS-Qwen 14B Mean
Full Model 4.46 5.38 3.10 4.31
GRIFFIN 1.61 7.20 5.00 4.60
Layer-wise Caprese 1.35 4.85 4.92 3.71
E2E Caprese 3.65 5.77 6.16 5.19
CATS 4.89 5.77 2.23 4.30
Layer-wise Caprese 3.22 5.90 4.80 3.97
E2E Caprese 2.64 4.95 5.28 4.29
@@ 4.1.1 Enhanced Performance with Reselection
For GRIFFIN, updating the metric $\bm{s}$ entails integrating the FF feature statistics of generated tokens into $\bm{s}$. While this can be done by re-running prefill on all tokens, there is a more efficient way by passing in the generated tokens following the last reselection through the full model. As these tokens propagate through each layer, we find the selection metric for the generated tokens $\bm{s}_{G}$ and update corresponding KV pairs. This is like verification in speculative decoding [7, 32]. As $\bm{s}$ and $\bm{s}_{G}$ are $\ell_{2}$-norms along the token axis, we define the updated metric as $\sqrt{(\bm{s}\odot\bm{s})+(\bm{s}_{G}\odot\bm{s}_{G})}$ and use that to reselect different subsets of the FF block to use (as described in Section 2). This updates the pruned layers yet avoids prefill for all tokens. Table 3 shows a clear benefit of reselection (even if infrequent compared to speculative decoding) in Caprese by pushing the performance much closer to the full model’s as we decrease steps between reselection rounds. Although GRIFFIN is generally more harmful to accuracy than CATS due to its structured pruning, with reselection, Caprese (GRIFFIN) can exceed the performance of Caprese (CATS) and reach the full model performance.
@@
Reselection is also possible with CATS but requires recomputing prefill for all tokens. The parameter to update is the hard thresholding parameter $\tau$, but since this requires finding the desired percentile of intermediate values in the FF block, we would need to access to all values, which likely cannot be fully stored in memory. In turn, prefill will need to be redone anytime we want to update $\tau$. Since $\tau$ represents a cutoff for a desired percentile (e.g., median), it is fairly robust to new observations. Given this and the lack of computational incentive, we focus primarily on reselection for GRIFFIN.
4.2 Scaling Best-of-$N$: Coverage
Now, we see the generalizability of Caprese on the second axis of test-time scaling: sampling multiple responses and selecting the best one, known as Best-of-$N$. Increasing the number of generations per prompt, $N$, increases the probability of generating a correct answer, even for non-thinking models (e.g., instruct models). To evaluate, we find the coverage (Pass@$K$) on 140 samples from MATH. We use an oracle verifier to accurately assess the quality of the pool of generated responses. To combat the high variance when $K$ is close to $N$, we calculate the average coverage for $K=1,\dots,100$ across 10 independent pools of 100 generations for each sample. Factoring in the saved compute, E2E Caprese is able to have similar or better coverage scaling compared to the original model, as shown in Figure 3. Notably, E2E Caprese (GRIFFIN) on Llama 3.2 3B Instruct improves Pass@$100$ by 7.0% from the original model while using 15.8% less compute.
Table 4: Instruct models’ 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER, XSum, and CNN/DailyMail). $F$ sparsity is set to 50%.
Model GSM8K MATH CoQA QASPER XSum CNN/DailyMail
Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43 21.65 25.60
GRIFFIN 7.13 5.42 56.05 14.11 21.13 25.47
Layer-wise Caprese 13.72 6.62 56.07 13.40 20.65 26.18
E2E Caprese 21.00 8.44 56.55 13.88 20.71 26.18
CATS 19.18 7.54 54.40 13.88 21.10 24.73
Layer-wise Caprese 18.65 8.28 55.58 14.35 20.94 25.35
E2E Caprese 20.39 9.04 56.12 13.75 20.40 25.56
Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45 23.22 26.20
GRIFFIN 28.96 10.98 64.52 12.52 22.09 25.49
Layer-wise Caprese 40.18 13.70 64.33 11.60 21.56 25.90
E2E Caprese 44.66 16.96 64.83 12.35 21.26 26.04
CATS 41.24 12.04 58.87 11.36 22.23 25.40
Layer-wise Caprese 45.49 13.62 60.53 12.26 22.50 25.90
E2E Caprese 46.85 14.40 61.72 12.36 21.58 25.40
Llama 3.1 8B Instruct 53.98 13.94 63.88 15.16 21.97 25.98
GRIFFIN 20.47 6.16 63.37 12.63 21.53 25.89
Layer-wise Caprese 37.60 9.72 65.05 14.21 21.92 26.31
E2E Caprese 51.40 12.64 65.50 15.35 22.05 26.42
CATS 50.95 11.80 58.85 12.26 21.43 25.67
Layer-wise Caprese 51.93 13.66 63.92 14.34 22.58 25.79
E2E Caprese 58.00 13.88 64.27 14.42 22.08 26.13
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96 22.17 26.01
GRIFFIN 33.74 11.32 63.28 11.07 18.27 22.24
Layer-wise Caprese 42.53 12.32 63.77 10.75 21.63 26.37
E2E Caprese 48.14 13.70 63.37 11.05 22.48 27.16
CATS 34.42 10.56 61.53 10.11 21.97 26.46
Layer-wise Caprese 46.32 13.90 61.92 10.82 22.10 26.23
E2E Caprese 46.17 14.16 63.92 11.03 22.15 26.52
Gemma 2 9B Instruct 78.17 27.64 63.78 9.91 23.98 26.46
GRIFFIN 59.21 25.22 63.82 10.14 24.66 26.77
Layer-wise Caprese 76.72 25.84 64.20 9.82 24.88 26.88
E2E Caprese 76.65 25.30 64.42 9.92 24.89 25.47
CATS 76.50 27.32 64.37 9.78 24.74 26.64
Layer-wise Caprese 77.18 28.16 64.52 10.46 24.54 26.45
E2E Caprese 77.18 28.00 64.87 10.02 24.65 26.33
4.3 Instruct Models
Next, we look into the performance of Caprese in the absence of inference scaling methods on instruct LLMs whcih are primarily tailored for language tasks. We show Caprese is able to preserve math performance without sacrificing quality on language tasks like question answering. We test Llama 3 [19] and Gemma 2 [51] models on 0-shot GSM8K [8], MATH [24], CoQA [43], QASPER [10], XSum [40], and CNN/DailyMail [25, 47]. We use CoT prompts for math tasks.
Table 4 shows Caprese preserves most if not all of the math capabilities in the original models without damaging performance on the language tasks, despite the distillation dataset being all math. In most cases, CATS and GRIFFIN severely harm GSM8K and MATH accuracy, but Caprese effectively recovers the lost performance. E2E Caprese is the most performative in the majority of math scenarios. Caprese’s performance is consistent at different sparsity levels and $r$ (Appendix A). All methods have little impact on the accuracy of these tasks (the main purpose of these tasks is to show no degradation in language-related cases).
@@ 4.4 Efficiency
Caprese reduces generation latency (Table 5). Caprese cuts total latency and time to next token by >16% for the Qwen 2.5 14B architecture. Moreover, the latency differences between GRIFFIN and Caprese are exceptionally small, suggesting that Caprese has minimal overhead. We also include latencies of GRIFFIN paired with LoRA weights, which although faster than the base model, negates >40% of the time savings that GRIFFIN has provided. This highlights the efficiency suboptimality of LoRA during inference. Metrics were collected on an NVIDIA L40 GPU using BF16 precision.
Table 5: End-to-end generation latency (s) and average time to next token (ms) for Qwen 2.5 14B. For the “Setup” column, $P+G$ indicates input and generation lengths of $P$ and $G$ tokens, respectively. As before, GRIFFIN is applied to the first half of the model, sparsity is 50%, and $r=256$.
@@ 4.4.1 Bonus: Effect on Natural Response Length
Caprese implicitly encourages brevity, often even producing shorter responses than from the full thinking model. Intriguingly, this behavior arises despite any enforcement or regularization on response lengths anywhere during training or inference. Shown in Figure 4, the shortest mean response length for all MATH-500 problem difficulties is either the full model or Caprese. CATS consistently outputs the longest responses, averaging $\sim$1K more across every sample. With increasing difficulty, all models and methods naturally allocate more inference tokens towards the output. Given CATS and Caprese (CATS) achieve similar MATH-500 accuracies to the full DS-Qwen 7B and 14B models, the ability of Caprese to cut down the lengthy responses of CATS down to the response lengths of the original model or shorter without compromising performance is a direct memory and latency benefit. This brevity of Caprese is more pronounced in larger models, proportionally, from a 5.8% reduction in tokens compared to the full DS-Qwen 7B model to an 8.5% reduction compared to the full DS-Qwen 14B model. See Appendix E for observations with GRIFFIN.
@@ 5 Conclusion
To combat the inefficiency of the long and brittle generation process associated with reasoning, we introduce Caprese, a performant and efficient method with strong reasoning and language capabilities that is compatible with a broad class of efficient FF algorithms. In the future, it would be interesting to evaluate its benefit in more reasoning domains, explore input-dependent low-rank layers, and investigate the knowledge contained in these low-rank layers. Caprese, along with future developments, pushes towards the long-term goal of degradation-free efficient LLM inference.
@@ Appendix C LoRA Performance
Table 6: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER, XSum, and CNN/DailyMail) with LoRA. $F$ sparsity is set to 50%, $r=256$, and LoRA ranks are set to match the number of parameters that Caprese adds.
Model GSM8K MATH CoQA QASPER XSum CNN/DailyMail
Llama 3.2 1B Instruct 22.44 10.66 55.43 14.43 21.65 25.60
GRIFFIN 7.13 5.42 56.05 14.11 21.13 25.47
Layer-wise LoRA 8.11 5.60 56.00 14.65 21.08 25.51
E2E LoRA 20.47 10.10 56.10 14.41 21.09 25.62
Layer-wise Caprese 13.72 6.62 56.07 13.40 20.65 26.18
E2E Caprese 21.00 8.44 56.55 13.88 20.71 26.18
Llama 3.2 3B Instruct 51.55 14.32 63.95 12.45 23.22 26.20
GRIFFIN 28.96 10.98 64.52 12.52 22.09 25.49
Layer-wise LoRA 26.46 11.28 63.88 12.55 22.12 25.96
E2E LoRA 42.91 16.50 64.67 12.63 21.81 26.05
Layer-wise Caprese 40.18 13.70 64.33 11.60 21.56 25.90
E2E Caprese 44.66 16.96 64.83 12.35 21.26 26.04
Gemma 2 2B Instruct 51.02 16.06 63.77 10.96 22.17 26.01
GRIFFIN 33.74 11.32 63.28 11.07 18.27 22.24
Layer-wise LoRA 39.58 14.66 62.98 10.54 21.58 26.64
E2E LoRA 44.66 16.71 63.03 10.81 22.45 26.35
Layer-wise Caprese 42.53 12.32 63.77 10.75 21.63 26.37
E2E Caprese 48.14 13.70 63.37 11.05 22.48 27.16
Appendix D Caprese with Quantization
To demonstrate recovery generalizability beyond sparse FF methods as the backend, we also equip QLoRA [13] with Caprese. Table 7 shows that E2E Caprese is also able to nearly recover the performance degraded by quantization in both math and language tasks.
Table 7: 0-shot accuracies on mathematical reasoning (GSM8K and MATH) and language generation tasks (CoQA, QASPER) with QLoRA as the backend method. $F$ weights are quantized to 4 bits and $r=256$.